# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v1

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly being used for odometry and SLAM applications. However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing point cloud matching based solutions, especially those originally intended for more accurate sensors such as LiDAR. Inspired by visual odometry research around 3D Gaussian Splatting [8], in this paper we propose using freely positioned 3D Gaussians to create a summarized representation of a radar point cloud tolerant to sensor noise, and subsequently leverage its inherent probability distribution function for registration (similar to NDT [2]). Moreover, we propose simultaneously optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function. Finally, we fuse our Gaussian modeling and scan matching algorithms into an EKF radar-inertial odometry system designed after current best practices. Experiments show that our Gaussian-based odometry is able to outperform current baselines on a well-known 4D radar dataset used for evaluation.
Our code and results can be publicly accessed at: https://github.com/robotics-upo/gaussian-rio
@@ I Introduction
Inspired by both of these lines of research, in this paper we propose and describe methods for obtaining a Gaussian model representation of a point cloud, and performing registration of a point cloud with respect to an existing Gaussian model. While this is similar in principle to the Normal Distributions Transform (NDT) [2], our methods consider Gaussians to be arbitrarily positioned in space instead of following a regular grid – like 3DGS [8]. Moreover, our scan matching (registration) step considers multiple simultaneous hypotheses, which can help increase the performance of the system, as we show in our experiments.
@@
Figure 2 shows the overall pipeline of our odometry system. As in other solutions, we use an Extended Kalman Filter (EKF) to perform inertial strapdown and process sensory input, including radar egovelocity. Keyframing is used to select individual 4D radar scans, which then undergo Gaussian modeling. The Gaussian multi-hypothesis scan matching algorithm registers every incoming 4D radar scan against the Gaussian model in order to produce a relative pose observation with respect to the last keyframe, which is then used to further update the EKF. Our experiments show how these Gaussian-based modeling and registration processes improve the results of existing state-of-the-art 4D radar-inertial odometry baselines.
@@ II Related work
II-A Inertial odometry and SLAM
There exist a large number of works focused on Inertial Odometry, sometimes even attaining a full SLAM pipeline (including loop closure). Particularly we are interested in approaches that use non-visual sensors, such as [23, 21, 15, 3, 4, 24, 25, 26, 22, 20]. These works use a 3D LiDAR (LIO) or 4D radar (RIO) sensor together with an IMU to provide an odometry solution. The main technical advantage of these sensors over visual sensors (cameras) is their increased field/depth of view and robustness to adverse environmental conditions: in particular, both are immune to low light conditions, and 4D radar in addition offers robustness to fog, dust and rain. Moreover, 4D radar is uniquely capable of measuring the velocity of objects by leveraging the Doppler effect. Researchers have taken advantage of this to directly measure the robot’s own velocity (“egovelocity”) [3, 4], something which has been shown to significantly improve the performance of odometry systems.
Registration of incoming point clouds with respect to previous ones is a key technique used in odometry systems. In particular, the most widely used algorithm for aligning point clouds is Generalized Iterative Closest Point (GICP) [17]. In the case of radar scans, this is a more difficult task compared to LiDAR because of the lower number and density of points, as well as higher measurement noise/uncertainty in each point’s position. For this reason [26] proposed the Adaptive Probability Distribution variant of the algorithm (APDGICP), which leverages GICP’s ability to make use of the spatial probability distribution of each point, calculating the required point position covariance matrices according to the radar specifications. However, this is still in essence an approach focused on individual points, without attempting to extract larger scale geometric features that may be more stable and robust for registration purposes.
@@ II-B Gaussian-based modeling
The Normal Distributions Transform (NDT) algorithm [2] slices and subdivides a point cloud using a regular grid and computes normal distributions for each cell that contains points. Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan. An important limitation of this algorithm is the fact that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment. Another problem lies in the fact that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization. Moreover, the original authors only consider 2D scan matching (requiring 3D point clouds to be projected into the XY plane), although subsequent works [11] extend the algorithm to 3D. NDT has been used with radar scans too [10], although only for pure radar odometry as opposed to radar-inertial odometry, and without leveraging radar egovelocity information.
3D Gaussian Splatting (3DGS) [8] is a popular technique for creating a model of a scene from multiple posed images. A 3DGS model is a collection of multiple 3D normal distributions (Gaussians) freely distributed in space, each also containing color/radiance information encoded using spherical harmonics. Even though this technique was originally developed as an alternative to NeRF [13] for novel view synthesis, researchers have successfully leveraged this representation to develop SLAM solutions [12, 7, 1]. However, 3DGS and these works are all designed for visual sensors (cameras), and they cannot leverage other sensors such as LiDAR or radar due to the nature of the representation. This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a SDF [14] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference.
@@ III-A Parametrization
  • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the natural logarithms of the square roots of the eigenvalues of the covariance matrix.
  • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the eigenvectors of the covariance matrix (see below); where $\mathbb{H}$ is the field of quaternions.
@@
In addition, the auxiliary parameters $\hat{\mathbf{s}}_{j}$ and $\hat{\mathbf{q}}_{j}$ are defined and used during optimization, such that ${\hat{\mathbf{s}}_{j}=\max_{\circ}(s_{\mathrm{min}},\mathbf{s}_{j})}$ and $\mathbf{q}_{j}=\frac{\hat{\mathbf{q}}_{j}}{|\hat{\mathbf{q}}_{j}|}$ ($\max_{\circ}$ being the element-wise max operator). This transformation allows us to define a minimum log scale factor $s_{\mathrm{min}}$ to prevent Gaussians from becoming infinitely small (and Mahalanobis distances infinitely big); and also enforce that $\mathbf{q}_{i}$ must be a unit quaternion. Typically we make the minimum size equal to the radar point uncertainty (standard deviation of position noise).
@@ III-B Initialization
The model $\mathcal{G}$ is initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\hat{\mathbf{q}}_{j}$ set to the unit quaternion. As for $\boldsymbol{\mu}_{j}$, we perform a Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
@@ III-C Optimization
We define the following loss function affecting a given $\theta_{j}$:
$\mathcal{L}_{j}=\frac{1}{2|G_{j}|}\sum_{\mathbf{p}_{i}\in G_{j}}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}+\sum_{k}^{3}\hat{\mathbf{s}}_{j_{k}}+\left(\min_{k}^{3}\hat{s}_{j\,k}-s_{\textrm{disc}}\right)^{+},$ (1)
where $x^{+}=\max(0,x)$, and $s_{\textrm{disc}}$ is a hyperparameter that introduces a log size prior for the smallest dimension of the Gaussian, encouraging it to be shaped like a thin disc (better suited to modeling surfaces).
To see how the other terms of the loss function are derived, first we introduce the log PDF of the multivariate normal distribution, which we need to maximize, so we flip its sign:
$-\log\mathrm{P}(\mathbf{p}_{i}|\theta_{j})=K+\frac{1}{2}\log|\boldsymbol{\Sigma}_{j}|+\\ +\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}\boldsymbol{\Sigma}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j}).$ (2)
[…]
$|\boldsymbol{\Sigma}_{j}|=|\mathbf{M}_{j}||\mathbf{M}_{j}^{T}|=|\mathbf{R}_{j}||\mathbf{S}_{j}|^{2}|\mathbf{R}_{j}^{T}|=|\mathbf{S}_{j}|^{2}=\left(\prod_{k}^{3}e^{\hat{s}_{j\,k}}\right)^{2},$ (3)
and thus,
$\frac{1}{2}\log|\boldsymbol{\Sigma}_{j}|=\frac{1}{2}2\sum_{k}^{3}\log e^{\hat{s}_{j\,k}}=\sum_{k}^{3}\hat{s}_{j\,k}.$ (4)
Likewise, we can simplify the term involving $p_{i}$. Note that it is half of the square of the Mahalanobis distance between the point and the Gaussian. The matrix $\boldsymbol{\Sigma}_{j}^{-1}$ can be decomposed as such: $\boldsymbol{\Sigma}_{j}^{-1}=(\mathbf{M}_{j}^{T})^{-1}\mathbf{M}_{j}^{-1}=(\mathbf{M}_{j}^{-1})^{T}\mathbf{M}_{j}^{-1}$. We can substitute it in the term and simplify:
$\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}\boldsymbol{\Sigma}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})=\frac{1}{2}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})^{T}(\mathbf{M}_{j}^{-1})^{T}\cdot\\ \cdot\mathbf{M}_{j}^{-1}(\mathbf{p}_{i}-\boldsymbol{\mu}_{j})=\frac{1}{2}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}.$ (5)
Lastly, we convert the (per-point) log probability function into a per-Gaussian loss function by taking the mean log probability value across all points assigned to the Gaussian. This has the effect of introducing the $\frac{1}{|G_{j}|}$ factor in the Mahalanobis term, and has no effect in the determinant term since it is constant for all points within the Gaussian. We decided to aggregate with the mean in order to ensure that every Gaussian equally contributes towards the loss function, instead of being biased towards Gaussians with larger $G_{j}$ sets. The resulting loss function $\mathcal{L}$ of the entire model $\mathcal{G}$ can be finally obtained by once again taking the mean of the loss values across all Gaussians.
@@ IV-B Optimization
The process is performed iteratively using gradient descent. Each epoch involves the following steps:
  • We define $K$ virtual copies of $\mathcal{P}$ called $\mathcal{P}^{\prime k}$, each registered according to every pose hypothesis $\hat{\xi}^{k}$. We also suggest downsampling the point cloud before creating the copies in order to reduce the computational load while retaining a representative sample of the geometry captured by the point cloud.
@@
  • We calculate the loss function and perform gradient descent on $\hat{\xi}^{k}$.
The loss function used is simply
$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min(d^{k}_{i},d_{\textrm{max}});\;\;\mathcal{L}=\sum_{i}^{K}\mathcal{L}_{i},$ (6)
where $d_{\textrm{max}}$ is the maximum Mahalanobis distance allowed, $\mathcal{L}_{k}$ is the loss function associated to a given particle $\hat{\xi}^{k}$, and $\mathcal{L}$ is the loss function affecting the entire swarm. Once the optimization process stops, we can select the particle with the lowest loss as the resulting output pose of the registration algorithm.
@@ V Inertial odometry with EKF
We use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results. Unlike other published solutions [18, 3, 4], we only use Error State filtering for the attitude, as opposed to the entire state. The state ($\mathbf{x}$), control ($\mathbf{u}$) and noise ($\mathbf{w}$) vectors of the filter are the following:
$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{v}\\ \mathbf{b}_{a}\\ \mathbf{b}_{\omega}\\ \delta\boldsymbol{\theta}\end{bmatrix};\;\mathbf{u}=\begin{bmatrix}\mathbf{a}\\ \boldsymbol{\omega}\end{bmatrix};\;\mathbf{w}=\begin{bmatrix}\mathbf{w}_{v}\\ \mathbf{w}_{\theta}\\ \mathbf{w}_{a}\\ \mathbf{w}_{\omega}\\ \mathbf{w}_{b_{a}}\\ \mathbf{w}_{b_{\omega}}\end{bmatrix},$ (7)
where $\mathbf{p}$ is the position, $\mathbf{v}$ is the velocity, $\mathbf{b}_{a}$ and $\mathbf{b}_{\omega}$ are the accelerometer/gyroscope biases, $\delta\boldsymbol{\theta}$ is the attitude error expressed as a tangential rotation $\in\mathfrak{so}(3)$, and $\mathbf{a}$ and $\boldsymbol{\omega}$ are the IMU accelerometer/gyroscope readings. Velocity and attitude process noise are modeled by $\mathbf{w}_{v}$ and $\mathbf{w}_{\theta}$ respectively. Note that the attitude quaternion $\mathbf{q}$ is not included in $\mathbf{x}$, and the attitude error $\delta\boldsymbol{\theta}$ (which takes its place in $\mathbf{x}$) is nominally always 0. We follow Kalibr’s IMU noise model [16] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.
We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$). The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$. We define $\mathbf{C}^{w}_{b}=\mathbf{R}\{\delta\mathbf{q}\}\mathbf{R}\{\mathbf{q}\}$, approximating $\mathbf{R}\{\delta\mathbf{q}\}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\approx\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization, and $\mathbf{C}^{b}_{r}$ is given as a parameter. All state parameters ($\mathbf{p}$, $\mathbf{v}$, $\mathbf{q}$) are expressed in the world frame.

### v2

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly being used for odometry and SLAM applications. However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing point cloud matching based solutions, especially those originally intended for more accurate sensors such as LiDAR. Inspired by visual odometry research around 3D Gaussian Splatting [10], in this paper we propose using freely positioned 3D Gaussians to create a summarized representation of a radar point cloud tolerant to sensor noise, and subsequently leverage its inherent probability distribution function for registration (similar to NDT [2]). Moreover, we propose simultaneously optimizing multiple scan matching hypotheses in order to further increase the robustness of the system against local optima of the function. Finally, we fuse our Gaussian modeling and scan matching algorithms into an EKF radar-inertial odometry system designed after current best practices. Experiments using publicly available 4D radar datasets show that our Gaussian-based odometry is comparable to existing registration algorithms, outperforming them in several sequences.
Our code and results can be publicly accessed at: https://github.com/robotics-upo/gaussian-rio-cpp
@@ I Introduction
In this paper we draw inspiration from both lines of research to propose the use of Gaussian models for representing a point cloud and performing registration with respect to it, as an alternative to existing registration methods such as the Normal Distributions Transform (NDT) [2] or Generalized Iterative Closest Point (GICP) [21]. Gaussians are freely positioned in space instead of following a regular grid, as in 3DGS and unlike NDT. Moreover, our scan matching (registration) step considers multiple simultaneous hypotheses, which can in some situations increase the performance of the system, as we show in our experiments. In addition, we outline an Extended Kalman Filter (EKF) based approach to radar-inertial odometry that consolidates existing practices by integrating both radar egovelocity [4] and keyframing/scan matching components, which we then use to evaluate our Gaussian-based approach in comparison to other registration algorithms on multiple datasets and radar types.
@@
Figure 2 shows the overall pipeline of our odometry system. As in other solutions, we use an Extended Kalman Filter (EKF) to perform inertial strapdown and process sensory input, including radar egovelocity. Keyframing is used to select individual 4D radar scans, which then undergo Gaussian modeling. The Gaussian multi-hypothesis scan matching algorithm registers every incoming 4D radar scan against the Gaussian model in order to produce a relative pose observation with respect to the last keyframe, which is then used to further update the EKF.
@@ II Related work
II-A Odometry and SLAM
There exist a large number of works focused on odometry, sometimes even attaining a full SLAM pipeline (including loop closure). Particularly we are interested in approaches that use non-visual sensors, such as [27, 25, 19, 4, 5, 28, 29, 30, 26, 24, 31]. These works use a 3D LiDAR or 4D radar sensor, sometimes together with an IMU, to provide an odometry solution. The main technical advantage of these sensors over visual sensors (cameras) is their increased field/depth of view and robustness to adverse environmental conditions: in particular, both are immune to low light conditions, and 4D radar in addition offers robustness to fog, dust and rain. Moreover, 4D radar is uniquely capable of measuring the velocity of objects by leveraging the Doppler effect. Researchers have taken advantage of this to directly measure the robot’s own velocity (“egovelocity”) [4, 5], something which has been shown to significantly improve the performance of odometry systems.
Registration of incoming point clouds with respect to previous ones is a key technique used in some of the cited odometry systems. In particular, the most widely used algorithm for aligning point clouds is Generalized Iterative Closest Point (GICP) [21]. This algorithm calculates distribution probability at each point based on their neighborhood, and optimizes a distribution-to-distribution scan matching function. In the case of radar scans, this is a more difficult task compared to LiDAR because of the lower number and density of points, as well as higher measurement noise/uncertainty in each point’s position. For this reason [30] proposed the Adaptive Probability Distribution variant of the algorithm (APDGICP), which leverages GICP’s ability to make use of the spatial probability distribution of each point, calculating the required point position covariance matrices according to the radar specifications. However, this is still in essence an approach focused on individual points, without attempting to extract larger scale geometric features that may be more stable and robust for registration purposes.
The works cited in this section use different sources of information, however many do not leverage all mentioned sources at the same time. Table I includes a list of surveyed methods.
@@ II-B Gaussian-based modeling
The Normal Distributions Transform (NDT) algorithm [2] slices and subdivides a point cloud using a regular grid and computes normal distributions for each cell that contains points. Scan matching is performed by maximizing the resulting probability distribution function with respect to an incoming scan. An important limitation of this algorithm is the fact that the grid structure itself imposes restrictions on the geometry, and the resulting model may not be fine grained enough to accurately model the geometry of the environment. Another problem lies in the fact that discontinuities in the probability function are produced when partially registered points cross cell boundaries during optimization. Moreover, the original authors only consider 2D scan matching (requiring 3D point clouds to be projected into the XY plane), although subsequent works [15] extend the algorithm to 3D. NDT has been used with radar scans too [13, 14], albeit without inertial information. Another similar algorithm is the Voxelized GICP (VGICP) [12], which also creates an NDT-like voxel-based representation to serve as the reference model, and combines it with the distribution-to-distribution based approach of GICP.
3D Gaussian Splatting (3DGS) [10] is a popular technique for creating a model of a scene from multiple posed images. A 3DGS model is a collection of multiple 3D normal distributions (Gaussians) freely distributed in space, each also containing color/radiance information encoded using spherical harmonics. Even though this technique was originally developed as an alternative to NeRF [17] for novel view synthesis, researchers have successfully leveraged this representation to develop SLAM solutions [16, 9, 1]. However, 3DGS and these works are all designed for visual sensors (cameras), and they cannot leverage other sensors such as LiDAR or radar due to the nature of the representation. This affects the rendering process in particular: while NeRF’s ray casting framework can be easily adapted to directly predict the output of a SDF [18] (therefore removing the volume rendering component from the system), the differentiable tile rasterizer is a core part of 3DGS that cannot be avoided, neither during training nor inference; thus requiring significant redesign efforts in order to synthesize other types of data such as LiDAR scans [3].
@@ III-A Parametrization
  • Scaling vector $\mathbf{s}_{j}\in\mathbb{R}^{3}$, indicating the natural logarithms of the square roots of the eigenvalues of the covariance matrix. We impose a minimum size restriction so that Gaussians do not become infinitely small (and Mahalanobis distances infinitely large).
  • Rotation quaternion $\mathbf{q}_{j}\in\mathbb{H}$, compactly encoding the eigenvectors of the covariance matrix (see below); where $\mathbb{H}$ is the field of quaternions. In order to ensure that the quaternions remain unitary during the optimization, we normalize them prior to usage.
@@ III-B Initialization
First, it is necessary to determine $N$, that is, the number of 3D Gaussians in the model. We propose choosing $N$ proportionally based on the number of points, so that richer scans can be represented with more Gaussians. The model $\mathcal{G}$ is then initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\mathbf{q}_{j}$ set to the unit quaternion. As for $\boldsymbol{\mu}_{j}$, we perform Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
@@ III-C Optimization
We define the following loss function affecting a given $\theta_{j}$, derived from the log PDF of the multivariate normal distribution:
$\mathcal{L}_{j}=\frac{1}{2|G_{j}|}\sum_{\mathbf{p}_{i}\in G_{j}}\hat{\mathbf{p}}_{i}^{T}\hat{\mathbf{p}}_{i}+\sum_{k}^{3}{\mathbf{s}_{j_{k}}.}$ (1)
The loss function $\mathcal{L}$ of the entire model $\mathcal{G}$ can be obtained by taking the mean of the loss values across all Gaussians.
@@ IV-B Optimization
The process is performed iteratively using Gauss-Newton. Each iteration involves the following steps:
  • We define $K$ virtual copies of $\mathcal{P}$ called $\mathcal{P}^{\prime k}$, each registered according to every pose hypothesis $\hat{\xi}^{k}$.
@@
  • We perform the usual Gauss-Newton update step for each particle $\hat{\xi}^{k}$, using the sums of squared Mahalanobis distances as the function to optimize, as in GICP. We also exclude points whose $d^{k}_{i}$ is greater than a given distance threshold $d_{\textrm{max}}$.
The optimization stops when all particles converge to a solution (i.e. when the update applied is under a given order of magnitude), or when a maximum number of iterations is reached; whichever happens first. The resulting output pose of the registration algorithm is the particle with the lowest value of the scoring function $\mathcal{L}_{k}$ defined as follows:
$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min(d^{k}_{i},d_{\textrm{max}}){,}$ (2)
where $M$ is the number of points in the scan. We consider a failure condition when the output particle has not converged.
@@ IV-C Keyframing
As a fallback, if no scans have returned a successful match during the past given time interval $t$, a keyframe is automatically generated.
@@ V Inertial odometry with EKF
We use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results. Unlike other published solutions [22, 4, 5], we only use Error State filtering for rotational variables, as opposed to the entire state. We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$). The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$. We decompose each $\mathbf{C}$ into an error part and a nominal quaternion part as follows: $\mathbf{C}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\mathbf{R}\{\mathbf{q}\}$, approximating the error part as $\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization.
The state ($\mathbf{x}$), control ($\mathbf{u}$) and noise ($\mathbf{w}$) vectors of the filter are the following:
$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{v}\\ {\mathbf{t}_{r}^{b}}\\ \mathbf{b}_{a}\\ \mathbf{b}_{\omega}\\ {\delta\boldsymbol{\theta}_{b}^{w}}\\ {\delta\boldsymbol{\theta}_{r}^{b}}\end{bmatrix};\;\mathbf{u}=\begin{bmatrix}\mathbf{a}\\ \boldsymbol{\omega}\end{bmatrix};\;\mathbf{w}=\begin{bmatrix}\mathbf{w}_{v}\\ \mathbf{w}_{\theta}\\ \mathbf{w}_{a}\\ \mathbf{w}_{\omega}\\ \mathbf{w}_{b_{a}}\\ \mathbf{w}_{b_{\omega}}\end{bmatrix},$ (3)
where $\mathbf{p}$ is the position, $\mathbf{v}$ is the velocity, $\mathbf{t}_{r}^{b}$ is the translation between radar and IMU, $\mathbf{b}_{a}$ and $\mathbf{b}_{\omega}$ are the accelerometer/gyroscope biases, each $\delta\boldsymbol{\theta}$ is the error part of its corresponding $\mathbf{C}$ frame expressed as a tangential rotation $\in\mathfrak{so}(3)$, and $\mathbf{a}$ and $\boldsymbol{\omega}$ are the IMU accelerometer/gyroscope readings. The position and velocity of the robot are expressed in the world frame, while the nominal rotations $\mathbf{q}_{b}^{w}$ and $\mathbf{q}_{r}^{b}$ are stored outside the filter state as quaternions. Velocity and attitude process noise are modeled by $\mathbf{w}_{v}$ and $\mathbf{w}_{\theta}$ respectively. We follow Kalibr’s IMU noise model [20] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.